OpenStudy (anonymous):

absolute value of 1-4x) -7 greater than or equal to -2

OpenStudy (anonymous):

solve the inequality for x

jimthompson5910 (jim_thompson5910):

|1-4x|-7 >= -2 |1-4x| >= -2+7 |1-4x| >= 5 1-4x >= 5 or 1-4x <= -5 -4x >= 5-1 or -4x <= -5-1 -4x >= 4 or -4x <= -6 x <= 4/(-4) or x >= -6/(-4) x<=-1 or x >= 3/2

OpenStudy (anonymous):

im sorry but it was suppose to be less than or equal to -2

OpenStudy (anonymous):

I solved it thank you

jimthompson5910 (jim_thompson5910):

|1-4x|-7 <= -2 |1-4x| <= -2+7 |1-4x| <= 5 1-4x <= 5 and 1-4x >= -5 1-4x >= -5 and 1-4x <= 5 -4x >= -5-1 and -4x <= 5-1 -4x >= -6 and -4x <= 4 x <= -6/(-4) and x >= 4/(-4) x <= 3/2 and x >= -1 x >= -1 and x <= 3/2 -1 <= x and x <= 3/2 -1 <= x <= 3/2

jimthompson5910 (jim_thompson5910):

If this is what you got, congrats

OpenStudy (anonymous):

it is thanks

jimthompson5910 (jim_thompson5910):

awesome